OpenStudy (anonymous):

Given cos A=-45 find angle A in degrees.Round answer to nearest hundredth

OpenStudy (jdoe0001):

\(\bf cos(\theta)={\color{red}{ \square }}\implies cos^{-1}[cos(\theta)]=cos^{-1}({\color{red}{ \square }})\implies \theta=cos^{-1}({\color{red}{ \square }})\)

OpenStudy (jdoe0001):

\(\bf cos(A)={\color{red}{ -45 }}\implies cos^{-1}[cos(a)]=cos^{-1}({\color{red}{ -45 }})\implies A=cos^{-1}({\color{red}{ -45 }})\)

OpenStudy (jdoe0001):

hmmm got a bit truncated... ok well \(\bf cos(A)={\color{red}{ -45 }}\implies cos^{-1}[cos(a)]=cos^{-1}({\color{red}{ -45 }})\\\quad\\\implies A=cos^{-1}({\color{red}{ -45 }})\)